\begin{thebibliography}{47}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Akbik et~al.(2018)Akbik, Blythe, and Vollgraf]{akbik2018contextual}
Alan Akbik, Duncan Blythe, and Roland Vollgraf.
\newblock Contextual string embeddings for sequence labeling.
\newblock In \emph{Proceedings of the 27th International Conference on
  Computational Linguistics}, pages 1638--1649, 2018.

\bibitem[Aleksander~Wawer(2018)]{poleval2018-ner}
Estera~Ma{\l}ek Aleksander~Wawer.
\newblock {Results of the PolEval 2018 Shared Task 2: Named Entity
  Recognition}.
\newblock \emph{Proceedings of the PolEval 2018 Workshop}, pages 53--62, 2018.

\bibitem[Antoun et~al.(2020)Antoun, Baly, and Hajj]{antoun2020arabert}
Wissam Antoun, Fady Baly, and Hazem Hajj.
\newblock Arabert: Transformer-based model for arabic language understanding.
\newblock \emph{arXiv preprint arXiv:2003.00104}, 2020.

\bibitem[Arkhipov et~al.(2019)Arkhipov, Trofimova, Kuratov, and
  Sorokin]{arkhipov-etal-2019-tuning}
Mikhail Arkhipov, Maria Trofimova, Yuri Kuratov, and Alexey Sorokin.
\newblock Tuning multilingual transformers for language-specific named entity
  recognition.
\newblock In \emph{Proceedings of the 7th Workshop on Balto-Slavic Natural
  Language Processing}, pages 89--93, Florence, Italy, August 2019. Association
  for Computational Linguistics.
\newblock \doi{10.18653/v1/W19-3712}.
\newblock URL \url{https://www.aclweb.org/anthology/W19-3712}.

\bibitem[Bojanowski et~al.(2017)Bojanowski, Grave, Joulin, and
  Mikolov]{bojanowski2017enriching}
Piotr Bojanowski, Edouard Grave, Armand Joulin, and Tomas Mikolov.
\newblock Enriching word vectors with subword information.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  5:\penalty0 135--146, 2017.

\bibitem[Cañete et~al.(2020)Cañete, Chaperon, Fuentes, and
  Pérez]{CaneteCFP2020}
José Cañete, Gabriel Chaperon, Rodrigo Fuentes, and Jorge Pérez.
\newblock Spanish pre-trained bert model and evaluation data.
\newblock In \emph{Practical ML for Developing Countries Workshop @ ICLR 2020},
  2020.

\bibitem[Conneau and Lample(2019)]{NIPS2019_8928}
Alexis Conneau and Guillaume Lample.
\newblock Cross-lingual language model pretraining.
\newblock In H.~Wallach, H.~Larochelle, A.~Beygelzimer, F.~d'~Alch\'{e}-Buc,
  E.~Fox, and R.~Garnett, editors, \emph{Advances in Neural Information
  Processing Systems 32}, pages 7059--7069. Curran Associates, Inc., 2019.
\newblock URL
  \url{http://papers.nips.cc/paper/8928-cross-lingual-language-model-pretraining.pdf}.

\bibitem[Conneau et~al.(2019)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzman, Grave, Ott, Zettlemoyer, and Stoyanov]{conneau2019unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzman, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov.
\newblock Unsupervised cross-lingual representation learning at scale.
\newblock \emph{arXiv preprint arXiv:1911.02116}, 2019.

\bibitem[Cui et~al.(2019)Cui, Che, Liu, Qin, Yang, Wang, and
  Hu]{chinese-bert-wwm}
Yiming Cui, Wanxiang Che, Ting Liu, Bing Qin, Ziqing Yang, Shijin Wang, and
  Guoping Hu.
\newblock Pre-training with whole word masking for chinese bert.
\newblock \emph{arXiv preprint arXiv:1906.08101}, 2019.

\bibitem[Czapla et~al.(2019)Czapla, Gugger, Howard, and
  Kardas]{czapla2019universal}
Piotr Czapla, Sylvain Gugger, Jeremy Howard, and Marcin Kardas.
\newblock Universal language model fine-tuning for polish hate speech
  detection.
\newblock \emph{Proceedings of the PolEval 2019 Workshop}, page 149, 2019.

\bibitem[Dadas(2019)]{dadas2019combining}
Slawomir Dadas.
\newblock Combining neural and knowledge-based approaches to named entity
  recognition in polish.
\newblock \emph{International Conference on Artificial Intelligence and Soft
  Computing}, pages 39--50, 2019.

\bibitem[Dadas et~al.(2020)Dadas, Pere{\l}kiewicz, and
  Po{\'s}wiata]{dadas-etal-2020-evaluation}
Slawomir Dadas, Micha{\l} Pere{\l}kiewicz, and Rafa{\l} Po{\'s}wiata.
\newblock {Evaluation of Sentence Representations in Polish}.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 1674--1680, Marseille, France, May 2020. European Language
  Resources Association.
\newblock ISBN 979-10-95546-34-4.
\newblock URL \url{https://www.aclweb.org/anthology/2020.lrec-1.207}.

\bibitem[Dai et~al.(2019)Dai, Yang, Yang, Carbonell, Le, and
  Salakhutdinov]{dai-etal-2019-transformer}
Zihang Dai, Zhilin Yang, Yiming Yang, Jaime Carbonell, Quoc Le, and Ruslan
  Salakhutdinov.
\newblock Transformer-{XL}: Attentive language models beyond a fixed-length
  context.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 2978--2988, Florence, Italy, July 2019.
  Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P19-1285}.
\newblock URL \url{https://www.aclweb.org/anthology/P19-1285}.

\bibitem[de~Vries et~al.(2019)de~Vries, van Cranenburgh, Bisazza, Caselli, van
  Noord, and Nissim]{de2019bertje}
Wietse de~Vries, Andreas van Cranenburgh, Arianna Bisazza, Tommaso Caselli,
  Gertjan van Noord, and Malvina Nissim.
\newblock Bertje: A dutch bert model.
\newblock \emph{arXiv preprint arXiv:1912.09582}, 2019.

\bibitem[Delobelle et~al.(2020)Delobelle, Winters, and
  Berendt]{delobelle2020robbert}
Pieter Delobelle, Thomas Winters, and Bettina Berendt.
\newblock Robbert: a dutch roberta-based language model.
\newblock \emph{arXiv preprint arXiv:2001.06286}, 2020.

\bibitem[Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
\newblock {BERT}: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota, June 2019. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/N19-1423}.
\newblock URL \url{https://www.aclweb.org/anthology/N19-1423}.

\bibitem[Heafield(2011)]{heafield2011kenlm}
Kenneth Heafield.
\newblock Kenlm: Faster and smaller language model queries.
\newblock \emph{Proceedings of the Sixth Workshop on Statistical Machine
  Translation}, pages 187--197, 2011.

\bibitem[Howard and Ruder(2018)]{howard2018universal}
Jeremy Howard and Sebastian Ruder.
\newblock Universal language model fine-tuning for text classification.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 328--339, 2018.

\bibitem[Jawahar et~al.(2019)Jawahar, Sagot, and Seddah]{jawahar:hal-02131630}
Ganesh Jawahar, Beno{\^i}t Sagot, and Djam{\'e} Seddah.
\newblock {What does BERT learn about the structure of language?}
\newblock In \emph{{ACL 2019 - 57th Annual Meeting of the Association for
  Computational Linguistics}}, Florence, Italy, July 2019.
\newblock URL \url{https://hal.inria.fr/hal-02131630}.

\bibitem[Kitaev et~al.(2020)Kitaev, Kaiser, and Levskaya]{kitaev2019reformer}
Nikita Kitaev, Lukasz Kaiser, and Anselm Levskaya.
\newblock Reformer: The efficient transformer.
\newblock In \emph{International Conference on Learning Representations}, 2020.

\bibitem[Koco{\'n} et~al.(2019)Koco{\'n}, Mi{\l}kowski, and
  Za{\'s}ko-Zieli{\'n}ska]{kocon-etal-2019-multi-level}
Jan Koco{\'n}, Piotr Mi{\l}kowski, and Monika Za{\'s}ko-Zieli{\'n}ska.
\newblock Multi-level sentiment analysis of {P}ol{E}mo 2.0: Extended corpus of
  multi-domain consumer reviews.
\newblock In \emph{Proceedings of the 23rd Conference on Computational Natural
  Language Learning (CoNLL)}, pages 980--991, Hong Kong, China, November 2019.
  Association for Computational Linguistics.
\newblock \doi{10.18653/v1/K19-1092}.
\newblock URL \url{https://www.aclweb.org/anthology/K19-1092}.

\bibitem[Kovaleva et~al.(2019)Kovaleva, Romanov, Rogers, and
  Rumshisky]{kovaleva-etal-2019-revealing}
Olga Kovaleva, Alexey Romanov, Anna Rogers, and Anna Rumshisky.
\newblock Revealing the dark secrets of {BERT}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 4365--4374, Hong Kong,
  China, November 2019. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/D19-1445}.
\newblock URL \url{https://www.aclweb.org/anthology/D19-1445}.

\bibitem[Kudo and Richardson(2018)]{kudo-richardson-2018-sentencepiece}
Taku Kudo and John Richardson.
\newblock {S}entence{P}iece: A simple and language independent subword
  tokenizer and detokenizer for neural text processing.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing: System Demonstrations}, pages 66--71, Brussels,
  Belgium, November 2018. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/D18-2012}.
\newblock URL \url{https://www.aclweb.org/anthology/D18-2012}.

\bibitem[Kuratov and Arkhipov(2019)]{kuratov2019adaptation}
Yuri Kuratov and Mikhail Arkhipov.
\newblock Adaptation of deep bidirectional multilingual transformers for
  russian language.
\newblock \emph{arXiv preprint arXiv:1905.07213}, 2019.

\bibitem[Lan et~al.(2020)Lan, Chen, Goodman, Gimpel, Sharma, and
  Soricut]{lan2019albert}
Zhenzhong Lan, Mingda Chen, Sebastian Goodman, Kevin Gimpel, Piyush Sharma, and
  Radu Soricut.
\newblock Albert: A lite bert for self-supervised learning of language
  representations.
\newblock In \emph{International Conference on Learning Representations}, 2020.

\bibitem[Le et~al.(2019)Le, Vial, Frej, Segonne, Coavoux, Lecouteux, Allauzen,
  Crabb{\'e}, Besacier, and Schwab]{le2019flaubert}
Hang Le, Lo{\"\i}c Vial, Jibril Frej, Vincent Segonne, Maximin Coavoux,
  Benjamin Lecouteux, Alexandre Allauzen, Beno{\^\i}t Crabb{\'e}, Laurent
  Besacier, and Didier Schwab.
\newblock Flaubert: Unsupervised language model pre-training for french.
\newblock \emph{arXiv preprint arXiv:1912.05372}, 2019.

\bibitem[Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{arXiv preprint arXiv:1907.11692}, 2019.

\bibitem[Marcińczuk et~al.(2013)Marcińczuk, Ptak, Radziszewski, and
  Piasecki]{MarPtaRadzPia:13}
Michał Marcińczuk, Marcin Ptak, Adam Radziszewski, and Maciej Piasecki.
\newblock {O}pen {D}ataset for {D}evelopment of {P}olish {Q}uestion {A}nswering
  {S}ystems.
\newblock In Z.~Vetulani and H.~Uszkoreit, editors, \emph{Proceedings of Human
  Language Technologies as a Challenge for Computer Science and
  Linguistics’13}, pages 479--483, Poznań, 2013. Fundacja UAM.

\bibitem[Marelli et~al.(2014)Marelli, Menini, Baroni, Bentivogli, Bernardi, and
  Zamparelli]{marelli-etal-2014-sick}
Marco Marelli, Stefano Menini, Marco Baroni, Luisa Bentivogli, Raffaella
  Bernardi, and Roberto Zamparelli.
\newblock A {SICK} cure for the evaluation of compositional distributional
  semantic models.
\newblock In \emph{Proceedings of the Ninth International Conference on
  Language Resources and Evaluation ({LREC}'14)}, pages 216--223, Reykjavik,
  Iceland, May 2014. European Language Resources Association (ELRA).
\newblock URL
  \url{http://www.lrec-conf.org/proceedings/lrec2014/pdf/363_Paper.pdf}.

\bibitem[Martin et~al.(2019)Martin, Muller, Su{\'a}rez, Dupont, Romary, de~la
  Clergerie, Seddah, and Sagot]{martin2019camembert}
Louis Martin, Benjamin Muller, Pedro Javier~Ortiz Su{\'a}rez, Yoann Dupont,
  Laurent Romary, {\'E}ric~Villemonte de~la Clergerie, Djam{\'e} Seddah, and
  Beno{\^\i}t Sagot.
\newblock Camembert: a tasty french language model.
\newblock \emph{arXiv preprint arXiv:1911.03894}, 2019.

\bibitem[Mikolov et~al.(2013)Mikolov, Sutskever, Chen, Corrado, and
  Dean]{mikolov2013distributed}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg~S Corrado, and Jeff Dean.
\newblock Distributed representations of words and phrases and their
  compositionality.
\newblock In \emph{Advances in neural information processing systems}, pages
  3111--3119, 2013.

\bibitem[Nguyen and Nguyen(2020)]{nguyen2020phobert}
Dat~Quoc Nguyen and Anh~Tuan Nguyen.
\newblock Phobert: Pre-trained language models for vietnamese.
\newblock \emph{arXiv preprint arXiv:2003.00744}, 2020.

\bibitem[Ogrodniczuk and Kopeć(2014)]{OGRODNICZUK14.1211}
Maciej Ogrodniczuk and Mateusz Kopeć.
\newblock The polish summaries corpus.
\newblock In Nicoletta Calzolari~(Conference Chair), Khalid Choukri, Thierry
  Declerck, Hrafn Loftsson, Bente Maegaard, Joseph Mariani, Asuncion Moreno,
  Jan Odijk, and Stelios Piperidis, editors, \emph{Proceedings of the Ninth
  International Conference on Language Resources and Evaluation (LREC'14)},
  Reykjavik, Iceland, may 2014. European Language Resources Association (ELRA).
\newblock ISBN 978-2-9517408-8-4.

\bibitem[Peng(2019)]{peng_2019}
Tony Peng.
\newblock The staggering cost of training sota ai models, Jun 2019.
\newblock URL
  \url{https://syncedreview.com/2019/06/27/the-staggering-cost-of-training-sota-ai-models/}.

\bibitem[Pennington et~al.(2014)Pennington, Socher, and
  Manning]{pennington2014glove}
Jeffrey Pennington, Richard Socher, and Christopher Manning.
\newblock Glove: Global vectors for word representation.
\newblock In \emph{Proceedings of the 2014 conference on empirical methods in
  natural language processing (EMNLP)}, pages 1532--1543, 2014.

\bibitem[Peters et~al.(2018)Peters, Neumann, Iyyer, Gardner, Clark, Lee, and
  Zettlemoyer]{peters2018deep}
Matthew Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark,
  Kenton Lee, and Luke Zettlemoyer.
\newblock Deep contextualized word representations.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, volume~1, pages 2227--2237, 2018.

\bibitem[Przepi{\'o}rkowski et~al.(2012)Przepi{\'o}rkowski, Banko, G{\'o}rski,
  and Lewandowska-Tomaszczyk]{przep-2012}
Adam Przepi{\'o}rkowski, Miros{\l}aw Banko, Rafa{\l}~L G{\'o}rski, and Barbara
  Lewandowska-Tomaszczyk.
\newblock {Narodowy Korpus Jezyka Polskiego [Eng.: National Corpus of Polish]}.
\newblock \emph{Wydawnictwo Naukowe PWN, Warsaw}, 2012.

\bibitem[Ptaszynski et~al.(2019)Ptaszynski, Pieciukiewicz, and
  Dyba{\l}a]{ptaszynski2019results}
Michal Ptaszynski, Agata Pieciukiewicz, and Pawe{\l} Dyba{\l}a.
\newblock {Results of the PolEval 2019 Shared Task 6: First Dataset and Open
  Shared Task for Automatic Cyberbullying Detection in Polish Twitter}.
\newblock \emph{Proceedings of the PolEval 2019 Workshop}, page~89, 2019.

\bibitem[Rybak et~al.(2020)Rybak, Mroczkowski, Tracz, and
  Gawlik]{rybak2020klej}
Piotr Rybak, Robert Mroczkowski, Janusz Tracz, and Ireneusz Gawlik.
\newblock {{KLEJ}: Comprehensive Benchmark for Polish Language Understanding}.
\newblock \emph{arXiv preprint arXiv:2005.00630}, 2020.

\bibitem[Shibuya and Hovy(2019)]{shibuya2019nested}
Takashi Shibuya and Eduard Hovy.
\newblock Nested named entity recognition via second-best sequence learning and
  decoding.
\newblock \emph{arXiv preprint arXiv:1909.02250}, 2019.

\bibitem[Souza et~al.(2019)Souza, Nogueira, and Lotufo]{souza2019portuguese}
Fabio Souza, Rodrigo Nogueira, and Roberto Lotufo.
\newblock Portuguese named entity recognition using bert-crf.
\newblock \emph{arXiv preprint arXiv:1909.10649}, 2019.
\newblock URL \url{http://arxiv.org/abs/1909.10649}.

\bibitem[Sun et~al.(2019)Sun, Wang, Li, Feng, Tian, Wu, and Wang]{sun2019ernie}
Yu~Sun, Shuohuan Wang, Yukun Li, Shikun Feng, Hao Tian, Hua Wu, and Haifeng
  Wang.
\newblock Ernie 2.0: A continual pre-training framework for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1907.12412}, 2019.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez,
  Kaiser, and Polosukhin]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock In \emph{Advances in neural information processing systems}, pages
  5998--6008, 2017.

\bibitem[Virtanen et~al.(2019)Virtanen, Kanerva, Ilo, Luoma, Luotolahti,
  Salakoski, Ginter, and Pyysalo]{virtanen2019multilingual}
Antti Virtanen, Jenna Kanerva, Rami Ilo, Jouni Luoma, Juhani Luotolahti, Tapio
  Salakoski, Filip Ginter, and Sampo Pyysalo.
\newblock Multilingual is not enough: Bert for finnish.
\newblock \emph{arXiv preprint arXiv:1912.07076}, 2019.

\bibitem[Wr{\'o}blewska and
  Krasnowska-Kiera{\'s}(2017)]{wroblewska-krasnowska-kieras-2017-polish}
Alina Wr{\'o}blewska and Katarzyna Krasnowska-Kiera{\'s}.
\newblock Polish evaluation dataset for compositional distributional semantics
  models.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 784--792,
  Vancouver, Canada, July 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1073}.
\newblock URL \url{https://www.aclweb.org/anthology/P17-1073}.

\bibitem[Xu et~al.(2020)Xu, Zhang, and Dong]{xu2020cluecorpus2020}
Liang Xu, Xuanwei Zhang, and Qianqian Dong.
\newblock {CLUECorpus2020}: A large-scale chinese corpus for
  pre-traininglanguage model.
\newblock \emph{arXiv preprint arXiv:2003.01355}, 2020.

\bibitem[Yang et~al.(2019)Yang, Dai, Yang, Carbonell, Salakhutdinov, and
  Le]{yang2019xlnet}
Zhilin Yang, Zihang Dai, Yiming Yang, Jaime Carbonell, Russ~R Salakhutdinov,
  and Quoc~V Le.
\newblock Xlnet: Generalized autoregressive pretraining for language
  understanding.
\newblock In \emph{Advances in neural information processing systems}, pages
  5754--5764, 2019.

\end{thebibliography}

\documentclass{article}

\usepackage{arxiv}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{lipsum}
\usepackage{multirow}
\usepackage{makecell}
\usepackage{tabularx}
\usepackage{graphicx}
\usepackage{caption}
\usepackage{enumitem}
\usepackage{textcomp}
\usepackage[page]{appendix}

\graphicspath{ {include/} }

\title{Pre-training Polish Transformer-based Language Models at Scale}

\author{
 Sławomir Dadas \\
  National Information Processing Institute\\
  Warsaw, Poland\\
  \texttt{sdadas@opi.org.pl} \\
   \And
 Michał Perełkiewicz \\
  National Information Processing Institute\\
  Warsaw, Poland\\
  \texttt{mperelkiewicz@opi.org.pl} \\
  \And
 Rafał Poświata \\
  National Information Processing Institute\\
  Warsaw, Poland\\
  \texttt{rposwiata@opi.org.pl} \\
}

\begin{document}
\maketitle

\begin{abstract}
Transformer-based language models are now widely used in Natural Language Processing (NLP). This statement is especially true for English language, in which many pre-trained models utilizing transformer-based architecture have been published in recent years. This has driven forward the state of the art for a variety of standard NLP tasks such as classification, regression, and sequence labeling, as well as text-to-text tasks, such as machine translation, question answering, or summarization. The situation have been different for low-resource languages, such as Polish, however. Although some transformer-based language models for Polish are available, none of them have come close to the scale, in terms of corpus size and the number of parameters, of the largest English-language models. In this study, we present two language models for Polish based on the popular BERT architecture. The larger model was trained on a dataset consisting of over 1 billion polish sentences, or 135GB of raw text. We describe our methodology for collecting the data, preparing the corpus, and pre-training the model. We then evaluate our models on thirteen Polish linguistic tasks, and demonstrate improvements over previous approaches in eleven of them.
\keywords{Language Modeling \and Natural Language Processing}
\end{abstract}

\section{Introduction}
Unsupervised pre-training for Natural Language Processing (NLP) has gained popularity in recent years. The goal of this approach is to train a model on a large corpus of unlabeled text, and then use the representations the model generates as an input for downstream linguistic tasks. The initial popularization of these methods was related to the successful applications of pre-trained word vectors (embeddings), the most notable of which include Word2Vec \citep{mikolov2013distributed}, GloVe \citep{pennington2014glove}, and FastText \citep{bojanowski2017enriching}. These representations have contributed greatly to the development of NLP. However, one of the main drawbacks of such tools was that the static word vectors did not encode contextual information. The problem was addressed in later studies by proposing context-dependent representations of words based on pre-trained neural language models. For this purpose, several language model architectures which utilize bidirectional long short-term memory (LSTM) layers have been introduced. The popular models such as ELMo \citep{peters2018deep}, ULMFiT \citep{howard2018universal}, and Flair \citep{akbik2018contextual}, have led to significant improvements in a wide variety of linguistic tasks. Shortly after, \citet{devlin-etal-2019-bert} introduced BERT - a different type of language model based on transformer \citep{vaswani2017attention} architecture. Instead of predicting the next word in a sequence, BERT is trained to reconstruct the original sentence from one in which some tokens have been replaced by a special \emph{mask token}. Since the text representations generated by BERT have proved to be effective for NLP problems - even those which were previously considered challenging, such as question answering or common sense reasoning - more focus has been put on transformer-based language models. As a result, in the last two years we have seen a number of new methods based on that idea, with some modifications in the architecture or the training objectives. The approaches that have gained wide recognition include RoBERTa \citep{liu2019roberta}, Transformer-XL \citep{dai-etal-2019-transformer}, XLNet \citep{yang2019xlnet}, Albert \citep{lan2019albert}, and Reformer \citep{kitaev2019reformer}.

The vast majority of research on both transformer-based language models and transfer learning for NLP is targeted toward the English language. This progress does not translate easily to other languages. In order to benefit from recent advancements, language-specific research communities must adapt and replicate studies conducted in English to their native languages. Unfortunately, the cost of training state-of-the-art language models is growing rapidly \citep{peng_2019}, which makes not only individual scientists, but also some research institutions unable to reproduce experiments in their own languages. Therefore, we believe that it is particularly important to share the results of research - especially pre-trained models, datasets, and source code of the experiments - for the benefit of the whole scientific community. In this article, we describe our methodology for training two language models for Polish language based on BERT architecture. The smaller model follows the hyperparameters of an English-language BERT-base model, and the larger version follows the BERT-large model. To the best of our knowledge, the latter is the largest language model for Polish available to date, both in terms of the number of parameters (355M) and the size of the training corpus (135GB). We have released both pre-trained models publicly\footnote{\url{https://github.com/sdadas/polish-roberta}}. We evaluate our models on several linguistic tasks in Polish, including nine from the KLEJ benchmark \citep{rybak2020klej}, and four additional tasks. The evaluation covers a set of typical NLP problems, such as binary and multi-class classification, textual entailment, semantic relatedness, ranking, and Named Entity Recognition (NER).

\subsection{Language-specific and multilingual transformer-based models}
In this section we provide an overview of models based on the transformer architecture for languages other than English. Apart from English, the language on which NLP research is most focused currently is Chinese. This is reflected in the number of pre-trained models available \citep{xu2020cluecorpus2020,chinese-bert-wwm,devlin-etal-2019-bert,sun2019ernie}. Other languages for which we found publicly available pre-trained models included: Arabic \citep{antoun2020arabert}, Dutch \citep{de2019bertje,delobelle2020robbert}, Finnish \citep{virtanen2019multilingual}, French \citep{martin2019camembert,le2019flaubert}, German, Greek, Italian, Japanese, Korean, Malaysian, Polish, Portuguese \citep{souza2019portuguese}, Russian \citep{kuratov2019adaptation}, Spanish \citep{CaneteCFP2020}, Swedish, Turkish, and Vietnamese \citep{nguyen2020phobert}. Models covering a few languages of the same family are also available, such as SlavicBERT (Bulgarian, Czech, Polish, and Russian) \citep{arkhipov-etal-2019-tuning} and NordicBERT\footnote{\url{https://github.com/botxo/nordic_bert}} (Danish, Norwegian, Swedish, and Finnish). The topic of massive multilingual models covering tens, or in some cases more than a hundred languages, has attracted more attention in recent years. The original BERT model \citep{devlin-etal-2019-bert} was released along with a multilingual version covering 104 languages. XLM \citep{NIPS2019_8928} (fifteen, seventeen and 100 languages) and XLM-R \citep{conneau2019unsupervised} (100 languages) were released in 2019. Although it was possible to use these models for languages in which no monolingual models were available, language-specific pre-training usually leads to better performance. To date, two BERT-base models have been made available for Polish: HerBERT \citep{rybak2020klej} and Polbert\footnote{\url{https://github.com/kldarek/polbert}}, both of which utilize BERT-base architecture.

\subsection{Contributions} 
Our contributions are as follows: 
1) We trained two transformer-based language models for Polish, consistent with the BERT-base and BERT-large architectures. To the best of our knowledge, the second model is the largest language model trained for Polish to date, both in terms of the number of parameters and the size of the training corpus. 
2) We proposed a method for collecting and pre-processing the data from the Common Crawl database to obtain clean, high-quality text corpora. 
3) We conducted a comprehensive evaluation of our models on thirteen Polish linguistic tasks, comparing them to other available transformer-based models, as well as recent state-of-the-art approaches.
4) We made the source code of our experiments available to the public, along with the pre-trained models. 

\section{Language model pre-training}
\label{sec:methodology}
In this section, we describe our methodology for collecting and pre-processing the data used for training BERT-base language models. We then present the details of the training, explaining our procedure and the selection of hyperparameters used in both models.

\subsection{Training corpus}
Transformer-based models are known for their high capacity \citep{jawahar:hal-02131630,kovaleva-etal-2019-revealing}, which means that they can benefit from large quantities of text. An important step in the process of creating a language model, therefore, is to collect a sufficiently large text corpus. We have taken into account that the quality of the text used for training will also affect the final performance of the model. The easiest way to collect a large language-specific corpus is to extract it from Common Crawl - a public web archive containing petabytes of data crawled from web pages. The difficulty with this approach is that web-based data is often noisy and unrepresentative of typical language use, which could eventually have a negative impact on the quality of the model. In response to this, we have developed a procedure for filtering and cleaning the Common Crawl data to obtain a high-quality web corpus. The procedure is as follows:
\begin{enumerate}[labelwidth=0pt, labelindent=0pt]
\item We download full HTML pages (WARC files in Common Crawl), and use the resulting metadata to filter the documents written in Polish language.
\item We use \emph{Newspaper3k}\footnote{\url{https://newspaper.readthedocs.io/en/latest/}} - a tool which implements a number of heuristics for extracting the \emph{main content} of the page, discarding any other text such as headers, footers, advertisements, menus, or user comments.
\item We then remove all texts shorter than 100 characters. Additionally, we identify documents containing the words: ‘przeglądarka’, ‘ciasteczka’, ‘cookies’, or ‘javascript’. The presence of these words may indicate that the extracted content is a description of a cookie policy, or default content for browsers without JavaScript enabled. We discard all such texts if they are shorter than 500 characters.
\item In the next step, we use a simple statistical language model (KenLM \citep{heafield2011kenlm}), trained on a small Polish language corpus to assess the quality of each extracted document. For each text, we compute the perplexity value and discard all texts with perplexity higher than 1000.
\item Finally, we remove all duplicated texts.
\end{enumerate}

The full training corpus we collected is approximately 135GB in size, and is composed of two components: the web part and the base part. For the web part, which amounts to 115GB of the corpus, we downloaded three monthly dumps of Common Crawl data, from November 2019 to January 2020, and followed the pre-processing steps described above. The base part, which comprises the remaining 20GB, is composed of publicly available Polish text corpora: the Polish language version of Wikipedia (1.5GB), the Polish Parliamentary Corpus (5GB), and a number of smaller corpora from the CLARIN (\url{http://clarin-pl.eu}) and OPUS (\url{http://opus.nlpl.eu}) projects, as well as Polish books and articles.

\subsection{Training procedure}

The authors of the original BERT paper \citep{devlin-etal-2019-bert} proposed two versions of their transformer-based language model: BERT-large (more parameters and higher computational cost), and BERT-base (fewer parameters, more computationally efficient). To train the models for Polish language, we adapted the same architectures. Let $L$ denote the number of encoder blocks, $H$ denote the hidden size of the token representation, and $A$ denote the number of attention heads. Specifically, we used $L=12,H=768, A=12$ for the base model, and $L=24,H=1024, A=16$ for the large model. The large model was trained on the full 135GB text corpus, and the base model on only the 20GB base part. The training procedure we employed is similar to the one suggested in the RoBERTa pre-training approach \citep{liu2019roberta}. Originally, BERT utilized two training objectives - Masked Language Modeling (MLM), and Next Sentence Prediction (NSP). We trained our models with the MLM objective, since it has been shown that NSP fails to improve the performance of the pre-trained models on downstream tasks \cite{liu2019roberta}. We also used dynamic token masking, and trained the model with a larger batch size than the original BERT. The base model was trained with a batch size of 8000 sequences for 125 000 training steps: the large model was trained with a batch size of 30 000 sequences for 50 000 steps. The reason for using such a large batch size for the bigger model is to stabilize the training process. During our experiments, we observed significant variations in training loss for smaller batch sizes, indicating that the initial combination of learning rate and batch size had caused an exploding gradient problem. To address the issue, we increased the batch size until the loss stabilized.

Both models were pre-trained with the Adam optimizer using the following optimization hyperparameters: $\epsilon = 1\mathrm{e}{-6},\beta_{1} = 0.9,\beta_{2} = 0.98$. We utilized a learning rate scheduler with linear decay. The learning rate is first increased for a warm-up phase of 10 000 update steps to reach a peak of $7\mathrm{e}{-4}$, and then linearly decreased for the remainder of the training. We also mimicked the dropout approach of the original BERT model: a dropout of 0.1 is applied on all layers and attention weights. The maximum length of a sequence was set to 512 tokens. We do not combine sentences from the training corpus: each is treated as a separate training sample. To encode input sequences into tokens, we employed SentencePiece \citep{kudo-richardson-2018-sentencepiece} Byte Pair Encoding (BPE) algorithm, and set the maximum vocabulary size to 50 000 tokens.

\section{Evaluation}
\label{sec:evaluation}
In this section, we discuss the process and results of evaluating our language models on thirteen Polish downstream tasks. Nine of these tasks constitute the recently developed KLEJ benchmark \citep{rybak2020klej}; three of them have already been introduced in \citet{dadas-etal-2020-evaluation}; and the last, named entity recognition, was a part of the PolEval\footnote{\url{http://2018.poleval.pl/index.php/tasks}} evaluation challenge. First, we compare the performance of our models with other Polish and multilingual language models evaluated on the KLEJ benchmark. Next, we present detailed per-task results, comparing our models with the previous state-of-the-art solutions for each of the tasks.

\subsection{Task descriptions}
\textbf{NKJP} 
(The National Corpus of Polish (Narodowy Korpus Języka Polskiego)) \cite{przep-2012} is one of the largest text corpora of the Polish language, consisting of texts from Polish books, news articles, web content, and transcriptions of spoken conversations. A part of the corpus, known as the ‘one million subcorpus’, contains annotations of named entities from six categories: ‘persName’, ‘orgName’, ‘geogName’, ‘placeName’, ‘date’, and ‘time’. The authors of the KLEJ benchmark used this subset to create a named entity classification task \cite{rybak2020klej}. First, they filtered out all sentences containing entities of more than one type. Next, they randomly assigned sentences to train development and test sets according to the rule that each named entity mentioned appears in only one of these splits. They undersample the ‘persName’ class, and merge the ‘date’ and ‘time’ classes to increase class balance.  Finally, they selected sentences without any named entity, and assigned them the ‘noEntity’ label. The resulting dataset consisted of 20 000 sentences belonging to six classes. The task is to predict the presence and type of each named entity. Classification accuracy is also reported.

\textbf{8TAGS} 
is a corpus created by \citet{dadas-etal-2020-evaluation} for their study on the subject of sentence representations in Polish language. This dataset was created automatically by extracting sentences from headlines and short descriptions of articles posted on the Polish social network, wykop.pl. It contains approximately 50 000 sentences, all longer than thirty characters, from eight popular categories: film, history, food, medicine, automotive, work, sport, and technology. The task is to assign a sentence to one of these classes in which classification accuracy is the measure.

\textbf{CBD} 
(Cyberbullying Detection) \cite{ptaszynski2019results} is a binary classification task, the goal of which is to determine whether a Twitter message constitutes a case of cyberbullying or not. This was a sub-task of task 6 in the PolEval 2019 competition. The dataset prepared by the competition’s organizers contains 11 041 tweets, extracted from nineteen of the most popular Polish Twitter accounts in 2017. The F1-score was used to measure the performance of the models.

\textbf{DYK}
‘Did you know?’ (\textit{‘Czy wiesz?’}) \cite{MarPtaRadzPia:13} is a dataset used for the evaluation and development of Polish language question answering systems. It consists of 4721 question-answer pairs obtained from the \textit{Czy wiesz...} Polish Wikipedia project. The answer to each question was found in the linked Wikipedia article. \citet{rybak2020klej} used this dataset to devise a binary classification task, the goal of which is to predict whether the answer to the given question is correct or not \cite{rybak2020klej}. Positive responses were additionally marked within larger fragments of responded text. Negative samples were selected by the BPE token overlap between a question and a possible answer. The F1-score was also reported for this task.

\textbf{PSC} 
The Polish Summaries Corpus \cite{OGRODNICZUK14.1211} is a corpus of manually created summaries of Polish language news articles. The dataset contains both abstract free-word summaries and extraction-based summaries created by selecting text spans from the original documents. Based on PSC, \cite{rybak2020klej} formulated a text-similarity task \cite{rybak2020klej}. They generate positive pairs by matching each extractive summary with the two least similar abstractive ones in the same article. Negative pairs were obtained by finding the two most similar abstractive summaries for each extractive summary, but from different articles. To calculate the similarity between summaries, they used the BPE token overlap. The F1-score was used for evaluation.

\textbf{PolEmo2.0} 
 \cite{kocon-etal-2019-multi-level} is a corpus of consumer reviews obtained from 
four domains: medicine, hotels, products, and school. Each of the reviews is annotated with one of four labels: positive, negative, neutral, or ambiguous. In general, the task is to choose the correct label, although here two special versions of the task are distinguished: PolEmo2.0-IN and PolEmo2.0-OUT. In PolEmo2.0-IN, both the training and test sets come from the same domains, namely medicine and hotels. In PolEmo2.0-OUT, however, the test set comes from the product and school domains. In both cases, accuracy was used for evaluation.

\textbf{Allegro Reviews} 
(AR) \cite{rybak2020klej} is a sentiment analysis dataset of product reviews from the e-commerce marketplace, allegro.pl. Each review has a rating on a five-point scale, in which one is negative, and five is positive. The task is to predict the rating of a given review. The macro-average of the mean absolute error per class (wMAE) is applied for evaluation.

\textbf{CDSC} 
(The Compositional Distributional Semantics Corpus) \cite{wroblewska-krasnowska-kieras-2017-polish} is a corpus of 10 000 human-annotated sentence pairs for semantic relatedness and entailment, in which image captions from forty-six thematic groups were used as sentences. Two tasks are proposed based on this dataset.
The CDSC-R problem involves predicting the relatedness between a pair of sentences, on a scale of zero to five, in which zero indicates that the sentences are not related, and five indicates that they are highly related. In this task, the Spearman correlation is used as an evaluation measure. CDSC-E’s task is to classify whether the premise entails the hypothesis (entailment), negates the hypothesis (contradiction), or is unrelated (neutral). For this task, accuracy is reported.

\textbf{SICK} 
 \cite{dadas-etal-2020-evaluation} is a manually translated Polish language version of the English
Natural Language Inference (NLI) corpus, SICK (Sentences Involving Compositional Knowledge) \cite{marelli-etal-2014-sick}, and consists of 10 000 sentence pairs. As with the CDSC dataset, two tasks can also be distinguished here. SICK-R is the task of predicting the probability distribution of relatedness scores (ranging from 1 to 5) for the sentence pair, in which the Spearman correlation is used for evaluation. SICK-E is a multiclass classification problem in which the relationship between two sentences is classified as entailment, contradiction, or neutral. Accuracy is used once again to measure performance.

\textbf{PolEval-NER 2018} 
 \cite{poleval2018-ner} was task 2 in the PolEval 2018 competition, the goal of which was to detect and assign the correct category and subcategory (if applicable) to a found named entity. In this study the task was simplified, as only the main categories had to be found. The effectiveness of the models is verified by the F1-score measure. This task was prepared on the basis of the NKJP dataset previously presented.

\subsection{Task-specific fine-tuning}

To evaluate our language models on downstream tasks, we fine-tuned them separately for each task. In our experiments, we encounter three types of problem: classification, regression, and Named Entity Recognition (NER). In classification tasks, the model is expected to predict a label from a set of two or more classes. Regression concerns the prediction of a continuous numerical value. NER is a special case of sequence tagging, i.e. predicting a label for each element in a sequence. The dataset for each problem consists of training and test parts, and in most cases also includes a validation part. The general fine-tuning procedure is as follows: we train our model on the training part of the dataset for a specific number of epochs. If the validation set is available, we compute the validation loss after each epoch, and select the model checkpoint with the best validation loss. For datasets without a validation set, we select the last epoch checkpoint. Then, we perform an evaluation on the test set using the selected checkpoint.

In the case of classification and regression tasks, we attach an additional fully-connected layer to the output of the \emph{[CLS]} token, which always remains in the first position of a sequence. For classification, the number of outputs for this layer is equal to the number of classes, and the softmax activation function is used. For regression, it is a linear layer with a single output. The models are fine-tuned with the Adam optimizer using the following hyperparameters: $\epsilon = 1\mathrm{e}{-6},\beta_{1} = 0.9,\beta_{2} = 0.98$. A learning rate scheduler with polynomial decay is utilized. The first 6\% of the training steps are reserved for the warm-up phase, in which the learning rate is gradually increased to reach a peak of $1\mathrm{e}{-5}$. By default, we train for ten epochs with a batch size of sixteen sequences. The specific fine-tuning steps and exceptions to the procedure are discussed below:
\begin{itemize}[leftmargin=*, labelwidth=0pt, labelindent=0pt]
\item \textbf{Classification on imbalanced datasets} – Some of the binary classification datasets considered in the evaluation, such as CBD, DYK, and PSC, are imbalanced, which means that they contain significantly fewer samples of the first class than of the second class. To counter this imbalance, we utilize a simple resampling technique: samples for the minority class in the training set are duplicated, and some samples for the majority class are randomly discarded. We set the resampling factor to 3 for the minority class, and 1 (DYK, PSC) or 0.75 (CBD) respectively for the majority class. Additionally, we increase the batch size for those tasks to thirty-two. 
\item \textbf{Regression} - In many cases, a regression task is restricted to a specific range of values for which the prediction is valid. For example, Allegro Reviews contains user reviews with ratings between one and five stars. For fine-tuning, we scale all the outputs of regression models to be within the range of $[0,1]$, and then rescale them to their original range during evaluation. Before rescaling, any negative prediction is set to 0, and any prediction greater that 1 is limited to 1. 
\item \textbf{Named entity recognition} - Since sequence tagging, in which the model is expected to generate per-token predictions, is different from simple classification or regression tasks, we decided to adapt an existing named entity recognition approach for fine-tuning using our language models. For this purpose, we employed a method from \citet{shibuya2019nested}, who proposed a transformer-based named entity recognition model with a Conditional Random Fields (CRF) inference layer, and multiple Viterbi-decoding steps to handle nested entities. In our experiments, we used the same hyperparameters as the authors.
\end{itemize}

\begin{table}
  \centering
  \caption{Results on the KLEJ benchmark.}
  \aboverulesep=0ex
  \belowrulesep=0ex
  \begin{tabular}{l|c|ccccccccc}
    \toprule
    \textbf{Model} & \textbf{Average} & NKJP & CDSC-E & CDSC-R & CBD & PE2-I & PE2-O & DYK & \makecell[{{p{0.7cm}}}]{\centering PSC} & \makecell[{{p{0.7cm}}}]{\centering AR} \\
    \hline
    \multicolumn{11}{l}{\textbf{Base models}}\\
    \hline
    \makecell[l]{mBERT} & 79.5 & 91.4 & 93.8 & 92.9 & 40.0 & 85.0 & 66.6 & 64.2 & 97.9 & 83.3 \\
    \makecell[l]{SlavicBERT} & 79.8	& 93.3 & 93.7 & 93.3 & 43.1 & 87.1 & 67.6 & 57.4 & 98.3 & 84.3 \\
    \makecell[l]{XLM-100} & 79.9 & 91.6 & 93.7 & 91.8 & 42.5 & 85.6 & 69.8 & 63.0 & 96.8 & 84.2 \\
    \makecell[l]{XLM-17} & 80.2 & 91.9 & 93.7 & 92.0 & 44.8 & 86.3 & 70.6 & 61.8 & 96.3 & 84.5 \\
    \makecell[l]{HerBERT} & 80.5 & 92.7 & 92.5 & 91.9 & 50.3 & 89.2 & 76.3 & 52.1 & 95.3 & 84.5 \\
    \makecell[l]{XLM-R base} & 81.5	& 92.1 & 94.1 & 93.3 & 51.0 & 89.5 & 74.7 & 55.8 & 98.2 & 85.2 \\
    \makecell[l]{Polbert} & 81.7 & 93.6 & 93.4 & 93.8 & 52.7 & 87.4 & 71.1 & 59.1 & 98.6 & 85.2 \\
    \textbf{Our model} & \textbf{85.3} & \textbf{93.9} & \textbf{94.2} & \textbf{94.0} & \textbf{66.7} & \textbf{90.6} & \textbf{76.3} & \textbf{65.9} & \textbf{98.8} & \textbf{87.8} \\
    \hline
    \multicolumn{11}{l}{\textbf{Large models}}\\
    \hline
    \makecell[l]{XLM-R large} & 87.5 & 94.1 & \textbf{94.4} & 94.7 & 70.6 & 92.4 & 81.0 & 72.8 & \textbf{98.9} & 88.4 \\
    \textbf{Our model} & \textbf{87.8} & \textbf{94.5} & 93.3 & \textbf{94.9} & \textbf{71.1} & \textbf{92.8} & \textbf{82.4} & \textbf{73.4} & 98.8 & \textbf{88.8} \\
    \bottomrule
  \end{tabular}
  \label{tab:eval_klej}
\end{table}

\subsection{Results and discussion}
In this section, we demonstrate the results of evaluating our language models on downstream tasks. We repeated the fine-tuning of the models for each task five times. The scores reported are the median values of those five runs. Table \ref{tab:eval_klej} demonstrates the evaluation results on the KLEJ benchmark, in comparison with other available Polish and multilingual transformer-based models. The results of other approaches are taken from the KLEJ leaderboard. We split the table into two sections, comparing the BERT-base and BERT-large architectures separately. We can observe that there is a wider selection of base models, and most of them are multilingual, such as the original multilingual BERT (mBERT) \citep{devlin-etal-2019-bert}, SlavicBERT \citep{arkhipov-etal-2019-tuning}, XLM \citep{NIPS2019_8928}, and XLM-R \citep{conneau2019unsupervised}. The only models pre-trained specifically for Polish language are HerBERT \citep{rybak2020klej} and Polbert. Among the base models, our approach outperforms others by a significant margin. In the case of large models, only the XLM-RoBERTa (XLM-R) pre-trained model has been available until now. XLM-RoBERTa is a recently published multilingual transformer trained on 2.5TB of data in 100 languages. It has been shown to be highly competitive against monolingual models. A direct comparison with our Polish language model demonstrates a consistent advantage of our model - it has achieved better results in seven of the nine tasks included in the KLEJ benchmark.

\begin{table}
  \centering
  \caption{Detailed results for Polish language downstream tasks. In some cases, we used the datasets and task definitions from the KLEJ benchmark, which are different from the original tasks they were based on (they have been reformulated or otherwise modified by the benchmark authors). We denote such tasks with (KLEJ) to emphasize that the evaluation was performed on the KLEJ version of the data. The abbreviated task types are: C - classification, R - regression, and ST - sequence tagging.}
  \aboverulesep=0ex
  \belowrulesep=0ex
  \begin{tabular}{lr|c|lc|c|c}
    \toprule
    \textbf{Task} & \makecell[{{p{0.7cm}}}]{~} & \textbf{Metric} & \textbf{Previous SOTA} & \makecell[{{p{0.9cm}}}]{~} & \makecell[c]{\textbf{Base}\\ \textbf{model}} & \makecell[c]{\textbf{Large}\\ \textbf{model}}\\
    \hline
    \multicolumn{5}{l}{\textbf{Multi-class classification}}\\
    \hline
    \makecell[l]{NKJP (KLEJ)} & \textbf{C} & Accuracy & XLM-R large \citep{conneau2019unsupervised} & 94.1 & 93.9 (\textminus0.2) & \textbf{94.5} (+0.4) \\
    \makecell[l]{8TAGS} & \textbf{C} & Accuracy & ELMo \citep{dadas-etal-2020-evaluation} & 71.4 & 77.2 (+5.8) & \textbf{80.8} (+9.4) \\
    \hline
    \multicolumn{5}{l}{\textbf{Binary classification}}\\
    \hline
    \makecell[l]{CBD} & \textbf{C} & F1-score & XLM-R large \citep{conneau2019unsupervised} & 70.6 & 66.7 (\textminus2.9) & \textbf{71.1} (+0.5)  \\
    \makecell[l]{DYK (KLEJ)} & \textbf{C} & F1-score & XLM-R large \citep{conneau2019unsupervised} & 72.8 & 65.9 (\textminus6.9) & \textbf{73.4} (+0.6) \\
    \makecell[l]{PSC (KLEJ)} & \textbf{C} & F1-score & XLM-R large \citep{conneau2019unsupervised} & \textbf{98.9} & 98.8 (\textminus0.1) & 98.8 (\textminus0.1) \\
    \hline
    \multicolumn{5}{l}{\textbf{Sentiment analysis}}\\
    \hline
    \makecell[l]{PolEmo2.0-IN} & \textbf{C} & Accuracy & XLM-R large \citep{conneau2019unsupervised} & 92.4 & 90.6 (\textminus1.8) & \textbf{92.8} (+0.4) \\
    \makecell[l]{PolEmo2.0-OUT} & \textbf{C} & Accuracy & XLM-R large \citep{conneau2019unsupervised} & 81.0 & 76.3 (\textminus4.7) & \textbf{82.4} (+1.4) \\
    \makecell[l]{Allegro Reviews} & \textbf{R} & 1-wMAE & XLM-R large \citep{conneau2019unsupervised} & 88.4 & 87.8 (\textminus1.0) & \textbf{88.8} (+0.4) \\
    \hline
    \multicolumn{5}{l}{\textbf{Textual entailment}}\\
    \hline
    \makecell[l]{CDSC-E} & \textbf{C} & Accuracy & XLM-R large \citep{conneau2019unsupervised} & \textbf{94.4} & 94.2 (\textminus0.2) & 93.3 (\textminus1.1)  \\
    \makecell[l]{SICK-E} & \textbf{C} & Accuracy & LASER \citep{dadas-etal-2020-evaluation} & 82.2 & 86.1 (+3.9) & \textbf{87.7} (+5.5) \\
    \hline
    \multicolumn{5}{l}{\textbf{Semantic relatedness}}\\
    \hline
    \makecell[l]{CDSC-R} & \textbf{R} & Spearman & XLM-R large \citep{conneau2019unsupervised} & 94.7 & 94.0 (\textminus0.7) & \textbf{94.9} (+0.2) \\
    \makecell[l]{SICK-R} & \textbf{R} & Spearman & USE \citep{dadas-etal-2020-evaluation} & 75.8 & 82.3 (+6.5) & \textbf{85.6} (+9.8) \\
    \hline
    \multicolumn{5}{l}{\textbf{Named entity recognition}}\\
    \hline
    \makecell[l]{Poleval-NER 2018} & \textbf{ST} & F1-score & \citet{dadas2019combining} & 86.2 & 87.9 (+1.7) & \textbf{90.0} (+3.8) \\
    \bottomrule
  \end{tabular}
  \label{tab:eval_tasks}
\end{table}

Table \ref{tab:eval_tasks} shows a more detailed breakdown of the evaluation results, and includes all the tasks from the KLEJ benchmark, and four additional tasks: SICK-R, SICK-R, 8TAGS, and PolEval-NER 2018. For each task, we define the task type (classification, regression, or sequence tagging), the metric used for evaluation, the previous state-of-the-art, and our results including the absolute difference to the SOTA. The competition between XLM-R and our large model dominates the results, since both models have led to significant improvements in linguistic tasks for Polish language. In some cases, the improvement over previous approaches is greater than 10\%. For example, the CDB task was a part of the PolEval 2019 competition, in which the winning solution by \citet{czapla2019universal} achieved an F1-score of 58.6. Both our model and the XLM-R large model outperform that by at least twelve points, achieving an F1-score of over 70. The comparison for the named entity recognition task is also interesting. The previous state-of-the-art solution by \citet{dadas2019combining} is a model that combined neural architecture with external knowledge sources, such as entity lexicons or a specialized entity linking module based on data from Wikipedia. Our language model managed to outperform this method by 3.8 points without using any structured external knowledge. In summary, our model has demonstrated an improvement over existing methods in eleven of the thirteen tasks. 

\section{Conclusions}
We have presented two transformer-based language models for Polish, pre-trained using a combination of publicly available text corpora and a large collection of methodically pre-processed web data. We have shown the effectiveness of our models by comparing them with other transformer-based approaches and recent state-of-the-art approaches. We conducted a comprehensive evaluation on a wide set of Polish linguistic tasks, including binary and multi-class classification, regression, and sequence labeling. In our experiments, the larger model performed better than other methods in eleven of the thirteen cases. To accelerate research on NLP for Polish language, we have released the pre-trained models publicly.

\bibliography{references}

\clearpage
\appendix
\clearpage
\section{Non-English transformer-based language models}

\begin{table*}[!htbp]
  \centering
  \begin{tabularx}{400pt}{llcll}
    \toprule
    \textbf{Project} & \textbf{Languages} & \textbf{Paper} & \textbf{Project URL} \\
    \midrule
Arabic-BERT & Arabic & - & \href{https://github.com/alisafaya/Arabic-BERT}{github.com/alisafaya/Arabic-BERT} \\
AraBERT & Arabic & \citep{antoun2020arabert} & \href{https://github.com/aub-mind/arabert}{github.com/aub-mind/arabert} \\
ClueCorpus2020 & Chinese & \citep{xu2020cluecorpus2020} & \href{https://github.com/CLUEbenchmark/CLUECorpus2020}{github.com/CLUEbenchmark/CLUECorpus2020} \\
Chinese BERT & Chinese & \citep{chinese-bert-wwm} & \href{https://github.com/ymcui/Chinese-BERT-wwm}{github.com/ymcui/Chinese-BERT-wwm} \\
Google BERT & Chinese & \citep{devlin-etal-2019-bert} & \href{https://github.com/google-research/bert}{github.com/google-research/bert} \\
ERNIE 2.0 & Chinese & \citep{sun2019ernie} & \href{https://github.com/PaddlePaddle/ERNIE}{github.com/PaddlePaddle/ERNIE} \\
BERTje & Dutch & \citep{de2019bertje} & \href{https://github.com/wietsedv/bertje}{github.com/wietsedv/bertje} \\
RoBBERT & Dutch & \citep{delobelle2020robbert} & \href{https://ipieter.github.io/blog/robbert}{ipieter.github.io/blog/robbert} \\
Finnish BERT & Finnish & \citep{virtanen2019multilingual} & \href{https://github.com/TurkuNLP/FinBERT}{github.com/TurkuNLP/FinBERT} \\
CamemBERT & French & \citep{martin2019camembert} & \href{https://camembert-model.fr}{camembert-model.fr} \\
FlauBERT & French & \citep{le2019flaubert} & \href{https://github.com/getalp/Flaubert}{github.com/getalp/Flaubert} \\
German BERT & German & - & \href{https://deepset.ai/german-bert}{deepset.ai/german-bert} \\
GreekBERT & Greek & - & \href{https://github.com/nlpaueb/greek-bert}{github.com/nlpaueb/greek-bert} \\
UmBERTo & Italian & - & \href{https://github.com/musixmatchresearch/umberto}{github.com/musixmatchresearch/umberto} \\
GilBERTo & Italian & - & \href{https://github.com/idb-ita/GilBERTo}{github.com/idb-ita/GilBERTo} \\
Japanese BERT & Japanese & - & \href{https://github.com/yoheikikuta/bert-japanese}{github.com/yoheikikuta/bert-japanese} \\
Japanese BERT & Japanese & - & \href{https://github.com/cl-tohoku/bert-japanese}{github.com/cl-tohoku/bert-japanese} \\
KoBERT & Korean & - & \href{https://github.com/SKTBrain/KoBERT}{github.com/SKTBrain/KoBERT} \\
Malaya & Malaysian & - & \href{https://github.com/huseinzol05/Malaya/}{github.com/huseinzol05/Malaya/} \\
Nordic BERT & Nordic (4) & - & \href{https://github.com/botxo/nordic\_bert}{github.com/botxo/nordic\_bert} \\
PolBERT & Polish & - & \href{https://github.com/kldarek/polbert}{github.com/kldarek/polbert} \\
HerBERT & Polish & \citep{rybak2020klej} & \href{https://klejbenchmark.com}{klejbenchmark.com} \\
Portuguese BERT & Portuguese & \citep{souza2019portuguese} & \href{https://github.com/neuralmind-ai/portuguese-bert}{github.com/neuralmind-ai/portuguese-bert} \\
RuBERT & Russian & \citep{kuratov2019adaptation} & \href{https://github.com/deepmipt/DeepPavlov}{github.com/deepmipt/DeepPavlov} \\
SlavicBERT & Slavic (4) & \citep{arkhipov-etal-2019-tuning} & \href{https://github.com/deepmipt/Slavic-BERT-NER}{github.com/deepmipt/Slavic-BERT-NER} \\
BETO & Spanish & \citep{CaneteCFP2020} & \href{https://github.com/dccuchile/beto}{github.com/dccuchile/beto} \\
Swedish BERT & Swedish & - & \href{https://github.com/Kungbib/swedish-bert-models}{github.com/Kungbib/swedish-bert-models} \\
BERTsson & Swedish & - & \href{https://huggingface.co/jannesg/bertsson}{huggingface.co/jannesg/bertsson} \\
BERTurk & Turkish & - & \href{https://github.com/stefan-it/turkish-bert}{github.com/stefan-it/turkish-bert} \\
PhoBERT & Vietnamese & \citep{nguyen2020phobert} & \href{https://github.com/VinAIResearch/PhoBERT}{github.com/VinAIResearch/PhoBERT} \\
    \bottomrule
  \end{tabularx}
  \label{tab:bert_models}
\end{table*}

\end{document}